OpenStudy (anonymous):

A 6 meter ladder is leaning on a vertical wall. The foot of the ladder is being pulled at a rate of 0.15m/s. Calculate the rate of which the height of the ladder is decreasing when the ladder is 3 meters off the ground.

OpenStudy (anonymous):

I started by doing \[\frac{ dH }{ dT } =\frac{ dH }{ dR }*\frac{ dR }{ dT }\]

OpenStudy (anonymous):

If that is correct, I need a formula for dH/dR which I don't have.

OpenStudy (anonymous):

 \[(x+0.15)^2+H^2=6^2\] I'm not sure, but, I think that's the one you're looking for

OpenStudy (anonymous):

solve for H...no wait, that's not right...unless you call the distance from the wall x, and then i'm kind of lost

OpenStudy (anonymous):

does that help at all?

OpenStudy (anonymous):

\[(0.15x)^2+H^2=6^2\]